OpenStudy (anonymous):

Solve the problem. Be sure to show all of your work. The wavelength, w, in meters per cycle, of a musical note is given by w= r/f where r is the speed of sound in meters per second and f is the frequency in cycles per second. The speed of sound in air is 344 m/sec. What is the wavelength of a note whose frequency in air is 33 cycles per second? Round to the nearest tenth of a meter per cycle.

sam (.sam.):

just divide w= r/f w= 344/33=10.42